0.2 mls of a 10-4 dilution of a virus preparation yields 90 plaques. What is the number of PFU per ml in the undiluted virus preparation
Answer #1

PFU (plaque forming unit) is the number of viral particles forming plaque per unit volume.

using the formula:

number of plaques formed / (dilution factor volume of diluted virus per well)

given: number of plaques formed = 90

number of dilutions = 10^4, this shows the sample is diluted 4 times in 10 ml of dilution per well.

dilution factor will be 1/10^4 = 0.0001

volume of diluted virus added = 0.2 ml

the number of plaques formed per unit in an undiluted sample =

pfu/ml

hence, the undiluted sample contains 4.5 106 pfu /ml.